3.2a a sample of propanol, C3H8O, contains 1.96 mol of the compound. Determine the amount (in...

3.2a

a sample of propanol, C3H8O, contains 1.96 mol of the compound. Determine the amount (in mol) of each element present and the number of atoms of each element present in the sample.

   Moles, Atoms

C

H

O

3.2b

Determine the percent composition of each element in manganese(II) phosphate, Mn3(PO4)2.

%Mn, %P, %O

3.2c

A compound is found to contain B, H, and Cl with the following percent composition:

B: 55.39%
H: 8.280%
Cl: 36.33%


What is the empirical formula of this compound?

Enter the elements
in the order B, H, Cl.
Answer #1

3.2. a) A sample of propanol (C3H8O) contains 1.96 moles of the compound. The molar masses of the elements are:

C: 12.0107 g/mol, H: 1.00794 g/mol, O: 15.999 g/mol.

The molar mass of propanol can be calculated as:

So the molar mass of propanol is 60.095 g/mol.

The percentage of each element present in the compound can be calculated as:

  

The mass of propanol in the compound is = 1.96 mol × 60.095 g/mol

       = 117.786 g.

So the number of moles of C   

  

Moles of H   

  

Moles of O

  

The number of atoms of each element is:

So for

C

H

O

b) Manganese phosphate has the formula Mn3(PO4)2 i.e. its empirical formula is Mn3P2O8.

The molar masses of the corresponding elements are:

Mn: 54.9380 g/mol, P: 30.9738 g/mol, O: 15.999 g/mol.

The molar mass of Mn3(PO4)2 is = (3×54.9380 g/mol) + (2×30.9738 g/mol) + (8×15.999 g/mol)

            = 354.7536 g/mol.

The percentage of each element is calculated as:

c) The percentages of the elements in a compound are:

                        B: 55.39 %, H: 8.280 %, Cl: 36.33%.

Assuming that the mass of the compound is 100 g, the masses of each of the elements can be determined as:

Mass of Element=Percentage of element×100 g.

Mass of B = (55.39/100) × 100 g = 55.39 g.

Mass of H = (8.280/100) × 100 g = 8.280 g.

Mass of Cl = (36.33/100) × 100 g = 36.33 g.

The molar masses of each element are:

B: 10.811 g/mol, H: 1.00794 g/mol, Cl: 35.453 g/mol.

From the masses, the number of moles of each element can be known.

Dividing each of the number of moles by the lowest number of moles i.e. dividing the number of moles of B, H, and Cl by the number of moles of Cl and rounding off to the nearest integer, we will get the molar ratio of the elements present.

So the empirical formula of the compound is B5H8Cl.